OpenStudy (anonymous):

A tank can be filled by one pipe in four hours and by a second pipe in six hours; and when it is full, the tank can be drained by a third pipe in three hours. If the tank is empty and all three pipes are open, in how many hours will the tank be filled? 8 hours 12 hours 24 hours are the the first and second pipe going at the same time ?????

OpenStudy (anonymous):

12

OpenStudy (anonymous):

can you awsner my question at the bottom

OpenStudy (anonymous):

No. Gimmie medal

OpenStudy (anonymous):

know you didnt help me in any way so no

OpenStudy (anonymous):

no*

OpenStudy (anonymous):

Both pipes are going st the same time.

OpenStudy (anonymous):

let v be the volume of tank,then in1 hr tank is full= \[\frac{ V }{ 4 }+\frac{ V }{ 6 }-\frac{ V }{ 3 }\] Calculate V

OpenStudy (anonymous):

how do i get v ?

OpenStudy (anonymous):

does one pipe fill up the tank in four hour while the other pipe goes for six hours .....do they fill up the tank as an addittive

OpenStudy (anonymous):

in 1 hr tank is full=\[V \frac{ 3+2-4 }{ 12 }=\frac{ V }{ 12 }\] OR Tank is full in hours=1*12=12hs

OpenStudy (anonymous):

we can say one 1/12 th part is full in hrs=1 tank is full in hrs=1*12/1=12 hrs.